# High Performance Distributed Control Law for Large-Scale Linear Systems: A Partitioned Distributed Observer Approach

### v3

@@ V-C Performance of the distributed-observer-based distributed control law
where $K_{ij}=\begin{bmatrix}\frac{k_{P_{i}}}{\tau_{P_{i}}}|\beta_{ij}|V_{i}V_{j}&0\end{bmatrix}$ and $K_{ii}$ is chosen so that $A_{ii}+B_{i}K_{ii}$ is a Hurwitz matrix. Subsequently, we obtain the distributed control law $\bar{u}_{i}$ by designing a partitioned distributed observer.
The observer gains $H_{i}$, and the weighted matrix $P_{i}$ can be calculated by Theorem 1 (Since there are $47$ different $H_{i}$ and $P_{i}$, we will not show their specific forms). We set $\gamma=100\theta^{2}$ and randomly select the initial values of the closed-loop system within $[0,1]$. In addition, the initial values of the partitioned distributed observer are all chosen as $\hat{x}_{ij}^{(p)}=2\times 1_{2|\mathcal{O}|_{p}}$ for all $\mathcal{O}_{p}\in\mathscr{P}_{i}$ and all $i=1,\ldots,47$.
Then, the trajectories of the closed-loop system with different $\theta$ are shown in Figure 11. Subfigure a) is the trajectories of $x_{c}(t)$ yielded by the centralized control law. Subfigures b), c), and d) are trajectories of $x_{r}(t)$, which is obtained by the distributed-observer-based distributed control law. From Figure 11d) and 11c), we know $x_{r}(t)$ cannot converge to zero when high-gain parameter $\theta=2$, while $x_{r}(t)$ tends to zero when $\theta=7$. However, it can be seen in Figure 11c) that the dynamic performance of $x_{r}(t)$ with $\theta=7$ is still not as good as that of $x_{c}(t)$. When $\theta=15$, we find that the trajectories in Figure 11b) are almost the same as that in Figure 11a). It indicates that the distributed-observer-based distributed control can achieve the same performance as that of the centralized control law.
@@
In addition, we need to point out that Figure 11 also shows the validity of Theorem 3. In this figure, closed-loop system states converge to a very small invariant set (The steady-state error shown in Figure 11b) is almost $5\times 10^{-6}$). In addition, the premise that the closed-loop system states can converge to any small invariant set is that the state estimation error of the partitioned distributed observer can converge to any invariant set at any fast speed. Therefore, the excellent performance of the closed-loop system in Figure 11b) infers the excellent performance of the partitioned distributed observer.
@@ VI Conclusions
This paper has developed a distributed control law based on the partitioned distributed observer for large-scale interconnected linear systems. Firstly, we have designed a partitioning algorithm, which is achieved by two steps, including initializing and merging. The algorithm can significantly reduce the dimension of local observers. Secondly, the partitioned distributed observer for large-scale systems has been designed. To analyze the stability of its error dynamics, this paper has proposed the two-layer Lyapunov analysis method and proved the dynamic transformation lemma of compact errors. Finally, we have designed the distributed control law based on the developed partitioned distributed observer, which has been proved to have the ability to approximate the control performance of the centralized control arbitrarily. The simulation results have shown that the relative error between the performance of the distributed control law and that of the centralized control law is less than $1\%$. Besides, the local observer dimension can be reduced by $90\%$ when the similarity between the communication network and the communication network is about $80\%$.

### v4

@@ V-C Performance of the distributed-observer-based distributed control law
where $K_{ij}=\begin{bmatrix}\frac{k_{P_{i}}}{\tau_{P_{i}}}|\beta_{ij}|V_{i}V_{j}&0\end{bmatrix}$ and $K_{ii}$ is chosen so that $A_{ii}+B_{i}K_{ii}$ is a Hurwitz matrix. Subsequently, we obtain the distributed control law $\bar{u}_{i}$ by designing a cover-based distributed observer.
The observer gains $H_{i}$, and the weighted matrix $P_{i}$ can be calculated by Theorem 1 (Since there are $47$ different $H_{i}$ and $P_{i}$, we will not show their specific forms). We set $\gamma=100\theta^{2}$ and randomly select the initial values of the closed-loop system within $[0,1]$. In addition, the initial values of the cover-based distributed observer are all chosen as $\hat{x}_{ij}^{(p)}=2\times 1_{2|\mathcal{O}|_{p}}$ for all $\mathcal{O}_{p}\in\mathscr{P}_{i}$ and all $i=1,\ldots,47$.
Then, the trajectories of the closed-loop system with different $\theta$ are shown in Figure 9. Subfigure a) is the trajectories of $x_{c}(t)$ yielded by the centralized control law. Subfigures b), c), and d) are trajectories of $x_{r}(t)$, which is obtained by the distributed-observer-based distributed control law. From Figure 9d) and 9c), we know $x_{r}(t)$ cannot converge to zero when high-gain parameter $\theta=2$, while $x_{r}(t)$ tends to zero when $\theta=7$. However, it can be seen in Figure 9c) that the dynamic performance of $x_{r}(t)$ with $\theta=7$ is still not as good as that of $x_{c}(t)$. When $\theta=15$, we find that the trajectories in Figure 9b) are almost the same as that in Figure 9a). It indicates that the distributed-observer-based distributed control can achieve the same performance as that of the centralized control law.
@@
In addition, we need to point out that Figure 9 also shows the validity of Theorem 3. In this figure, closed-loop system states converge to a very small invariant set (The steady-state error shown in Figure 9b) is almost $5\times 10^{-6}$). In addition, the premise that the closed-loop system states can converge to any small invariant set is that the state estimation error of the cover-based distributed observer can converge to any invariant set at any fast speed. Therefore, the excellent performance of the closed-loop system in Figure 9b) infers the excellent performance of the cover-based distributed observer.
@@ VI Conclusions
This paper has developed a distributed control law based on the cover-based distributed observer for large-scale interconnected linear systems. Firstly, we have designed a coverage solving algorithm, which is achieved by two steps, including initializing and merging. The algorithm can significantly reduce the dimension of local observers. Secondly, the cover-based distributed observer for large-scale systems has been designed. To analyze the stability of its error dynamics, this paper has proposed the two-layer Lyapunov analysis method and proved the dynamic transformation lemma of compact errors. Finally, we have designed the distributed control law based on the developed cover-based distributed observer, which has been proved to have the ability to approximate the control performance of the centralized control arbitrarily. The simulation results have verified the effectiveness of our theories.